you are the office manager of r a small company. You need to acquire a new copier for the office. You find a suitable one that leases for $250 a month from the copy machine company. It costs 2.5 cents per copy to run the machine. You purchases paper for $3.50 a ream (500 sheets). If your copying budget is no more than $950 per month, is this machine a good choice? Write a brief recommendation to the purchasing department. Use equations and inequalitities to explain your recommendation.
Answer by checkley71(8403)

950=250+X(3.5/500+.025)
SOLVING FOR X WHICH IS THE MAX COPYING FOR THE BUDGET
950=250+X(.007+.025)
950-250=.032X
.032X=700
X=700/.032
X=21,875 SHEETS CAN BE COPIED PER MONTH WITHIN THE BUDGET.
############################################################
HAPPY NEW YEAR
###########################################################
THE ONLY RECOMMENDATION YOU CAN MAKE IS IF THEY NEED TO COPY LESS THAN 21,875 SHEETS PER MONTH THAN THIS MACHINE WILL KEEP THEM BELOW THE BUDGET.